[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Aug 23, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Aug 24 02:22:10 UTC 2017


HSU Shift Log for Wed Aug 23 2017 22:22:04 GMT-0400 (EDT)

Shift: R

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: GOES16 meso request active through 9Z. Midnight shift  
should renew it (remember we are priority 4 now, not 24).
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170824/069eb89c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list